A beam of length 7m and mass 200kg is suspended from two ropes as show. Find the tension in each rope.

A beam of length 7m and mass 200kg is suspended fr
Answer #1

ANSWER

Weight of the beam
200 kg(9.8 m/s²) = 1960 N

Let T1 and T2 be the rope tensions.

forces in the vertical direction must add to zero
T1cos30 + T2cos60 - 1960 = 0

forces in the horizontal direction must add to zero
T1sin30 - T2sin60 = 0
T1 = T2sin60/sin30

substitute into the first equation

T2sin60cos30/sin30 + T2cos60 - 1960 = 0
T2(sin60cos30/sin20 + cos60) = 1960
T2 = 727.85 N

T1 = T2sin60/sin30
T1 = 727.85sin60/sin30
T1 = 1260.67 N

Regards!!!